OpenStudy (anonymous):

Use identities to find the exact value of the four remaining trig functions: Sin=1/2, cos=square root 3/2 I have that the tan is square root 3/3 Can someone help me find the other four functions?

OpenStudy (campbell_st):

ok... this is an exact value triangle  you need cosec(x) = 1/sin(x) so take the reciprocal of sin(x) sec(x) = 1/cos(x) so take the reciprocal of cos(x) cot(x) = 1/tan(x) so take the recprocal of tan(x) an alternate method is to use cosec(x) = hyp/opp sec(x) = adj/hyp cot(x) = adj/opp hope it helps

OpenStudy (anonymous):

So if I'm getting cot first, it would be tan flipped, then multiple top and bottom by square root 3?

OpenStudy (anonymous):

@campbell_st